OpenStudy (anonymous):

A rectangular prism has a length of 4", a width of 8", and a height of 12". If both the length and height double, how many times larger is the new volume than the original volume?

OpenStudy (campbell_st):

find the volume of the given prism \[V_{0} = 4 \times 8 \times 12 \] now double the measurements of length , now 8" and height now 24" and calculate the volume \[V_{1} = 8 \times 8 \times 24\] then the comparison will be \[\frac{V_{1}}{V_{0}} = \]